Debonairs Pizza discovered that their Pizza delivery times are normally distributed with a mean of 45 minutes and standard deviation of 8 minutes. What is the probability that: a) The Pizza they delivered last took between 45 minutes and 50 minutes to deliver b) The next Pizza will take less than 47 minutes to deliver


1
Expert's answer
2021-05-02T14:07:07-0400

"\mu = 45 \\

\sigma = 8"

a)

"P(45<X<50) = P(\frac{45-45}{8}<Z<\frac{50-45}{8}) \\

= P(0<Z<0.625) \\

= P(Z<0.625) -P(Z<0) \\

= 0.7340 -0.5 \\

= 0.2340"

b)

"P(X<47) = P(Z< \frac{47-45}{8}) \\

= P(Z<0.25) \\

= 0.5987"